Ok so i need to find x and y.
-2u+v=2
4u-4v=4

Answer by colliefan(242)

This system of equations has u and v so we'll solve for u and v.
-2u+v=2
4u-4v=4

Let's solve this by the substitution method. We will pick one of the equations and solve it for one of the variables. Then, we will take that expression that represents that variable and substitute it into the other equation so that we have one equation with one variable and we can solve that equation to get the value of one variable and then use that to calculate the value of the other variable.
Step by step:
Pick one equation and rearrange it to get an expression for one of the variables.
-2u+v=2
-2u+v+2u=2+2u
-2u+2u+v=2+2u
0+v=2+2u
v=2+2u
Now, substitute that expression into the other equation and solve that for the one variable it now contains.
4u-4v=4
4u-4(2+2u)=4
4u-8-8u=4
-8-4u=4
-8-4u+8=4+8
-4u=12
(-1/4)(-4)u=12*(-1/4)
1*u=-3
u=-3
Now, use that value of one variable and put it into either of the original equations to calculate the value of the other variable.
-2u+v=2
-2(-3)+v=2
6+v=2
v=-4
Check to see if the values for the variables are valid solutions to both of the original equations.